# BLADE: Box-Level Supervised Amodal Segmentation through Directed Expansion

### 04 Jan 2024 v2

@@ Introduction
BBTP (Hsu et al. 2019) Yes 66.5 59.7 58.4 57.9 54.4 51.0 48.9 50.4 44.7 40.2 53.2
BoxInst (Tian et al. 2021) Yes 72.3 52.5 53.5 53.9 37.7 38.1 38.2 23.0 22.8 23.7 41.6
Bayesian-Amodal (Sun, Kortylewski, and Yuille 2022) Yes 63.9 59.7 59.6 59.7 57.2 56.8 56.8 55.0 53.9 53.4 57.6
Bayesian-Amodal (Sun, Kortylewski, and Yuille 2022) No 63.0 59.5 59.5 59.5 56.2 55.9 55.6 51.9 50.6 48.3 56.0
@@
SAM (ViT-H) (Kirillov et al. 2023) - Yes 86.7 75.0 50.8 39.0 62.9 82.7 74.9 59.2 42.3 64.8
VRSP (Xiao et al. 2021) fully - 84.7 75.8 74.5 67.1 75.5 82.1 77.7 74.5 72.9 76.8
AISFormer (Tran et al. 2022) fully - 85.8 76.4 75.0 69.4 76.7 80.6 76.9 70.9 62.1 72.6
BBTP (Hsu et al. 2019) weakly Yes 77.0 68.3 58.9 53.9 64.5 57.3 49.4 40.7 35.0 45.6
BoxInst (Tian et al. 2021) weakly Yes 82.0 73.3 56.6 43.6 63.9 76.8 67.0 57.2 34.0 58.8
Bayesian-Amodal (Sun, Kortylewski, and Yuille 2022) weakly Yes 72.3 69.6 66.2 58.5 66.7 65.3 65.0 64.3 61.4 64.0
Bayesian-Amodal (Sun, Kortylewski, and Yuille 2022) weakly No 69.9 68.1 63.2 47.3 62.1 58.3 59.8 58.6 53.5 57.6
@@
Table 2: The ablation study on the KINS dataset. UN, NE, and FS respectively represent the uniform loss, the neighbor loss, and the fusion structure. The fusion structure indicates whether to adopt $L^{r}$ and the corresponding region-branch. All these experiments are conducted with unknown $c$.
@@
1.0 1.0 1.0 81.3 72.3 69.6 62.1 71.3
2.0 1.0 1.0 81.6 74.5 73.7 63.6 73.4
1.0 2.0 1.0 82.0 73.4 72.7 64.8 73.2
1.0 1.0 2.0 79.9 71.6 68.3 60.1 70.0
Table 2: The ablation study on the KINS dataset. UN, NE, and FS respectively represent the uniform loss, the neighbor loss, and the fusion structure. The fusion structure indicates whether to adopt $L^{r}$ and the corresponding region-branch. All these experiments are conducted with unknown $c$.
Table 3: The comparison of different weights in $L^{a}$ on the KINS dataset. All other experimental settings are the same.
Comparison with Existing Methods
Baselines.
Synthetic Occlusion.
Real Occlusion.
As shown in Table 1, our method also performs well on the KINS dataset which is based on real-world images. Although our method is weakly-supervised, which is not comparable to fully-supervised methods, we provide some results in Table 1 for better evaluation. Among all these methods, our approach achieves very competitive mean performance. Compared with Bayesian-Amodal (unknown $c$), our model improves by over 11% in mean IoU.
Transferability.
Ablation Study
The Effect of Fusion Structure.
To verify the effectiveness of the fusion structure, we cancel the region-branch and the structure of fusing multiple branches’ results in the baselines at the last four rows, directly taking the prediction of the amodal-branch as output. Compared with these experimental results, the corresponding results with the fusion structure show significantly better performance.
The Effect of Neighbor Loss.
The Effect of Uniform Loss.
The Effect of Different Weights in $L^{a}$.
As shown in Table 3, small adjustments of the weights in $L^{a}$ will result in certain but not dramatic performance changes, and our currently selected weights $\alpha_{1}^{a}=2.0,\alpha_{2}^{a}=1.0,\alpha_{3}^{a}=1.0$ achieve good performance.
Conclusion
In this work, we studied the problem of box-level supervised amodal segmentation. Noticing the issue of low resolution and inaccurate boundaries in existing methods which cannot meet the needs of practical applications, we propose a novel approach through directed expansion. Our proposed method introduces a structure of multi-branch fusion based on the overlapping region - the area where different instances intersect. Conservative strategy and expansion-encouraged strategy are applied to non-overlapping regions and overlapping regions, respectively. Utilizing predicted visible masks as clues, a connectivity loss is incorporated for reasonable expansion towards corresponding amodal masks. Consequently, experimental results indicate our method significantly outperforms other state-of-the-art methods, especially in images with high occlusion ratios.

### 25 Feb 2024 v3

@@ Introduction
BBTP Yes 66.5 59.7 58.4 57.9 54.4 51.0 48.9 50.4 44.7 40.2 53.2
BoxInst Yes 72.3 52.5 53.5 53.9 37.7 38.1 38.2 23.0 22.8 23.7 41.6
Bayesian-Amodal Yes 63.9 59.7 59.6 59.7 57.2 56.8 56.8 55.0 53.9 53.4 57.6
Bayesian-Amodal No 63.0 59.5 59.5 59.5 56.2 55.9 55.6 51.9 50.6 48.3 56.0
@@
SAM (ViT-H) - Yes 86.7 75.0 50.8 39.0 62.9 82.7 74.9 59.2 42.3 64.8
VRSP fully - 84.7 75.8 74.5 67.1 75.5 82.1 77.7 74.5 72.9 76.8
AISFormer fully - 85.8 76.4 75.0 69.4 76.7 80.6 76.9 70.9 62.1 72.6
BBTP weakly Yes 77.0 68.3 58.9 53.9 64.5 57.3 49.4 40.7 35.0 45.6
BoxInst weakly Yes 82.0 73.3 56.6 43.6 63.9 76.8 67.0 57.2 34.0 58.8
Bayesian-Amodal weakly Yes 72.3 69.6 66.2 58.5 66.7 65.3 65.0 64.3 61.4 64.0
Bayesian-Amodal weakly No 69.9 68.1 63.2 47.3 62.1 58.3 59.8 58.6 53.5 57.6
@@
Comparison with Existing Methods
Baselines.
Synthetic Occlusion.
Real Occlusion.
As shown in Table 1, our method also performs well on the KINS dataset which is based on real-world images. Although our method is weakly-supervised, which is not comparable to fully-supervised methods, we provide some results in Table 1 for better evaluation including SAM (Kirillov et al. 2023), VRSP (Xiao et al. 2021), and AISFormer (Tran et al. 2022). Among all these methods, our approach achieves very competitive mean performance. Compared with Bayesian-Amodal (unknown $c$), our model improves by over 11% in mean IoU.
Transferability.
Ablation Study
The Effect of Fusion Structure.
To verify the effectiveness of the fusion structure, we cancel the region-branch and the structure of fusing multiple branches’ results in the baselines from the 4th row to the 8th row, directly taking the prediction of the amodal-branch as output. Compared with these experimental results, the corresponding results with the fusion structure show significantly better performance.
The Effect of Neighbor Loss.
The Effect of Uniform Loss.
The Effect of Different Weights in $L^{a}$.
Small adjustments of the weights in $L^{a}$ will result in certain but not dramatic performance changes, and our selected weights $\alpha_{1}^{a}=2.0,\alpha_{2}^{a}=1.0,\alpha_{3}^{a}=1.0$ achieve good performance.
Conclusion
In this work, we achieve box-level supervised amodal segmentation through directed expansion. Our approach introduces a fusion structure based on the overlapping region. Conservative strategy and expansion-encouraged strategy are applied to non-overlapping regions and overlapping regions, respectively. Utilizing predicted visible masks as clues, a connectivity loss is incorporated for reasonable expansion. Experimental results indicate our method significantly outperforms other state-of-the-art methods.
@@ Conclusion
a 1.0 1.0 1.0 81.3 72.3 69.6 62.1 71.3
b 2.0 1.0 1.0 81.6 74.5 73.7 63.6 73.4
c 1.0 2.0 1.0 82.0 73.4 72.7 64.8 73.2
d 1.0 1.0 2.0 79.9 71.6 68.3 60.1 70.0
Table 2: The ablation study on the KINS dataset. UN, NE, and FS respectively represent the uniform loss, the neighbor loss, and the fusion structure. The fusion structure indicates whether to adopt $L^{r}$ and the corresponding region-branch.